OpenStudy (anonymous):

determine the sum or difference in clock 12 arithmetic. (7+10) + (8+7) A.7 B.8 C.32 D.2

jimthompson5910 (jim_thompson5910):

add up (7+10) + (8+7) to get ???

OpenStudy (anonymous):

32?

jimthompson5910 (jim_thompson5910):

good, now divide that into 12 32/12 = ??? remainder ???

OpenStudy (anonymous):

2.6 ? so the answer is 2...?

jimthompson5910 (jim_thompson5910):

Similar example: 40/12 = 3 remainder 4 This is because 40 goes into 12 three whole times and has 4 left over

jimthompson5910 (jim_thompson5910):

good so far

jimthompson5910 (jim_thompson5910):

but you're not quite done yet

jimthompson5910 (jim_thompson5910):

take that decimal part 0.6 and multiply it by 12 to get the remainder

jimthompson5910 (jim_thompson5910):

it should be 0.6667 btw

OpenStudy (anonymous):

it was. & i got 7.2 !

jimthompson5910 (jim_thompson5910):

it should be more like 0.667*12 = 8.004, which is close to 8 So 32/12 = 2 remainder 8

OpenStudy (anonymous):

ok.. so my answer is B.8 ...?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

Since "mod" deals with remainders

OpenStudy (anonymous):

thankyou!!

jimthompson5910 (jim_thompson5910):

you're welcome